# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

### v2

@@ Abstract
Here, we investigate the stability of steady-states of Eqs. (5) in the Letter, in combination with the aforementioned approximation techniques, using linear stability analysis. This allows us to identify the region in parameter space where we have a so-called Hopf bifurcation, which marks the transition from a (non-oscillatory) steady-state to coherent oscillations.
S4.1 The “monomer” approximation
Since in the MF approximation the local order is slaved by the global order, it suffices to consider the linear stability analysis for the global order. Let us consider a small perturbation of the global order around the steady-state value $\textbf{m}_{\rm s}$, i.e., $\mathbf{m}(t)=\mathbf{m}_{\rm s}+\mathbf{\delta m}(t)$. The most interesting steady-state value to consider is given by the disordered state $\mathbf{m}_{\rm s}=0$, in which case the probability (S27)-(S28) can be expanded as
@@
Inserting this linearized expression back into Eq. (S19), we eventually obtain
@@
Note that $\hat{M}_{\mu\mu}$ can be written in terms of the steady-state solution for the local order given by Eq. (S20), i.e., $\hat{M}_{\mu\mu}=zq^{\mu\mu}_{s}-1$, which we will make use of in the remaining calculation. Furthermore, the off-diagonals are related to the steady-state of the local order between the lattices, i.e., $\hat{M}_{ab}+\hat{M}_{ba}=2q^{ab}_{s}$. Hence, a perturbation of the global order couples to the steady-state value of the local order. The solution of Eq. (S47) can be written in terms of an eigenmode expansion
@@
Since all matrix entries in (S47) are real, the characteristic polynomial also has real coefficients. Therefore, if the eigenvalues are complex, they come in complex-conjugate pairs. The perturbation $\mathbf{\delta m}(t)$ grows in time when ${\rm Re}(\hat{\lambda}_{\pm})>0$, and shrinks in time when ${\rm Re}(\hat{\lambda}_{\pm})<0$. The imaginary part of the eigenvalues tells us whether the perturbation develops oscillations in time ${\rm Im}(\hat{\lambda}_{\pm})\neq 0$, or is monotonic in time ${\rm Im}(\hat{\lambda}_{\pm})=0$. The Hopf bifurcation, also known as an oscillatory instability or type-${\rm II}_{\rm o}$ instability in the Cross-Hohenberg classification [3], occurs when the complex conjugate eigenvalues cross the imaginary axis in the complex plane. Based on Eq. (S52), this occurs when
@@
S4.1.1 Perfectly nonreciprocal setting
@@
where the latter equation is directly satisfied for $K\neq 0$ (a proof is given in S4.2.2.). Hence, also in the “monomer” expression we find a critical value for the local order, which is given by $q_{\rm crit}\equiv 1/z$. Upon inserting Eq. (S48) into Eq. (S58), one can obtain an explicit result for the Hopf line in the $(J,K)-$plane.
S4.2 The Bethe-Guggenheim “pair” approximation
@@
For the sake of simplicity, we directly focus on the perfectly nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$.
S4.2.1 Steady-state local order
The trivial steady state is given by the disordered state with $m^{\mu}_{s}=0$ and $q^{ab}_{s}=0$. To solve for the steady state of the local order, denoted as $q^{aa}_{s}=q^{bb}_{s}\equiv q_{s}(J,K)$, we need to solve
$q_{s}(J,K)=\frac{1}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{a}_{l,n})},$ (S62)
where $\overline{\mathcal{P}}^{\pm}_{l}$ are the probabilities (S43)-(S44) evaluated at steady-state values
$\overline{\mathcal{P}}^{\pm}_{l}\equiv\mathcal{C}^{z}_{l}(1\mp q_{s})^{z-l}(1\pm q_{s})^{l}.$ (S63)
Equation (S62) can be solved for specific integer values of $z$. For example, for $z\in\{2,3\}$ we obtain
$q_{s}(J,K)|_{z=2}$ $=\frac{2-\sqrt{4-\left[\sum_{n=\pm}\tanh{(2J+nK)}\right]^{2}}}{\sum_{n=\pm}\tanh{(2J+nK)}},$ (S64)
$q_{s}(J,K)|_{z=3}$ $=\frac{4{-}\sqrt{16{-}\sum_{n=\pm}[\tanh{(J{+}nK)}{-}3\tanh{(3J{+}nK)}]\sum_{n=\pm}[\tanh{(J{+}nK)}{+}\tanh{(3J{+}nK)}]}}{\sum_{n=\pm}[\tanh{(J{+}nK)}{-}3\tanh{(3J{+}nK)}]}.$ (S65)
For $z=4$ the solution can be compactly written as the solution to a quartic equation
$q_{s}(J,K)|_{z=4}=\mathcal{S}(J,K)-(1/2)\sqrt{-4\mathcal{S}(J,K)^{2}+2\mathcal{H}(J,K)+\mathcal{Q}(J,K)/\mathcal{S}(J,K)},$ (S66)
where we have introduced the auxiliary functions
$\mathcal{H}(J,K)$ $\equiv\frac{3\cosh{(4J)}[\cosh{(4J)}+\cosh{(2K)}]}{\sinh^{2}{(2J)}[\cosh{(4J)}-2\sinh^{2}{(K)}]},$ (S67)
$\mathcal{Q}(J,K)$ $\equiv 16\left(\sum\nolimits_{n=\pm}[\tanh{(4J+nK)}-2\tanh{(2J+nK)}]\right)^{-1},$ (S68)
$\mathcal{S}(J,K)$ $\equiv(1/2)\sqrt{(2/3)\mathcal{H}(J,K)+(\mathcal{Q}(J,K)/6)\left(\mathcal{A}(J,K)+\Delta_{0}(J,K)/\mathcal{A}(J,K)\right)},$ (S69)
$\mathcal{A}(J,K)$ $\equiv 2^{-1/3}\left(\Delta_{1}(J,K)+\sqrt{\Delta^{2}_{1}(J,K)-4\Delta^{3}_{0}(J,K)}\right)^{1/3},$ (S70)
$\Delta_{0}(J,K)$ $\equiv-(3/4)\sum\nolimits_{n=\pm}[\tanh{(2J{+}nK)}+\tanh{(4J{+}nK)}]\sum\nolimits_{n=\pm}[\tanh{(2J{+}nK)}-\tanh{(4J{+}nK)}],$ (S71)
$\Delta_{1}(J,K)$ $\equiv\!216\cosh{(2J)}\!\cosh{(2K)}\!\operatorname{sech}{(4J{-}K)}\!\operatorname{sech}{(4J{+}K)}\!\sinh^{3}{(2J)}\sinh^{2}{(K)}[\cosh{(4J)}{+}\cosh{(2K)}]^{-2}.$ (S72)
Upon inserting the steady-state values, the probability can be expanded up to first order in $\delta\mathbf{m}(t)$ and $\delta\mathbf{q}(t)$, which reduces Eqs. (5) in the main Letter to the following linear set of equations
@@
where $\mathbf{M}(J,K)$ is a $2\times 2$ matrix and $\mathbf{Q}(J,K)$ a $3\times 3$ matrix. Due to the diagonal block structure of the linearized equations, we can handle the perturbations for $\delta\mathbf{m}(t)$ and $\delta\mathbf{q}(t)$ separately.
S4.2.2 Linear stability of global order
The elements of the $2\times 2$ matrix $\mathbf{M}$ are given by
$M_{aa}(J,K)$ $=M_{bb}(J,K)=\frac{q_{s}/q_{\rm crit}-1}{1+q_{s}}[1-2\sum_{l=0}^{z}\sum_{n=0}^{1}\overline{\mathcal{P}}^{+}_{l}\tanh(U^{a}_{l,n})],$ (S74)
$M_{ab}(J,K)$ $={-}M_{ba}(J,K)=\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\tanh(U^{a}_{l,n}),$ (S75)
where $q_{\rm crit}\equiv 1/(z-1)$. Note that Eqs. (S74)-(S75) are equivalent to Eqs. (7) in the Letter. Further analysis regarding the solution for $\delta\mathbf{m}(t)$ is given in the Letter. Here, we are left to prove that $M_{ab}(J,K\neq 0)\neq 0$. To see this, let us explicitly write out the first sum over $n$
$M_{ab}=\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\tanh(U^{a}_{l,n})=\sum_{l=0}^{z}(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\left(\tanh{([2l{-}z]J{+}K)}{-}\tanh{([2l{-}z]J{-}K)}\right),$ (S76)
Note that $\overline{\mathcal{P}}^{\pm}_{l}\geq 0$ for $q_{s}\in[-1,1]$, which follows straightforwardly from Eq. (S63). Furthermore, since $\tanh{(x)}$ is an increasing function of $x$, we have $\tanh{([2l-z]J+K)}-\tanh{([2l-z]J-K)}>0$ for $K>0$ and $\tanh{([2l-z]J+K)}-\tanh{([2l-z]J-K)}<0$ for $K<0$. Hence, $M_{ab}$ is given by a sum over strictly positive (for $K>0$) or negative (for $K<0$) terms, rendering $M_{ab}\neq 0$ for $K\neq 0$. This outcome results in complex eigenvalues for $\mathbf{M}$, as explained in the Letter.
S4.2.3 Linear stability of local order
The elements of the $3\times 3$ matrix $\mathbf{Q}$ are given by
@@
where for the last inequality we again used the triangle inequality. This establishes that $Q_{11}\leq 0$ and $Q_{33}\leq 0$, and therefore when $\tilde{\lambda}_{2,3}$ become complex, their real part obeys the bound ${\rm Re}(\tilde{\lambda}_{2,3})=(Q_{11}+Q_{33})/2\leq 0$. Hence, up to first order, any perturbation $\delta\mathbf{q}(t)$ decays over time. The coherent oscillations in $\mathbf{q}(t)$ observed in Fig. 1c,d in the main Letter are therefore an inherently nonlinear effect related to the coupling between $\mathbf{m}(t)$ and $\mathbf{q}(t)$.
Appendix S5 Divergence of the residence time upon approaching the SNIPER bifurcation
Consider the perfectly nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$. Figure 2c in the main Letter shows the magnetization oscillating between $4$ ghost states upon approaching the degenerate saddle-node-infinite-period (SNIPER) bifurcation from below, i.e. $J\uparrow J_{\rm SNP}(K)$, where $J_{\rm SNP}(K)$ is the $J$ value of the SNIPER bifurcation for a given nonreciprocal coupling $K$. The residence time $\mathcal{T}$ in the vicinity of these ghost states diverges at the SNIPER bifurcation as (see also Eq. (9) in the main Letter)
$\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2},$ (S91)
which is shown explicitly in Fig. S1. To see the precise algebraic scaling, we plot $1/\mathcal{T}^{2}$ against $J_{\rm SNP}(K)-J$ in Fig. S1b, where we indeed observe very good agreement with the expected scaling (black dashed line) close to the SNIPER bifurcation.
Figure S1: Algebraic divergence of the residence time $\mathcal{T}$ within a ghost state close to the degenerate saddle-node-infinite-period (SNIPER) bifurcation. Approaching the SNIPER bifurcation from below, the residence time scales as $\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2}$, where $J_{\rm SNP}(K)$ is the $J$ value of the SNIPER bifurcation for a given nonreciprocal coupling $K$. In both panels we consider $K=0.1$ and $z=4$, for which $J_{\rm SNP}\approx 0.4477$. Results are obtained with the numerical continuation package MATCONT [4].
@@
S6.1 Recap of the global order
@@
where $\{\boldsymbol{\sigma}\}$ denotes the set of all possible spin configurations $\boldsymbol{\sigma}$. From the master equation (S2) we can obtain the time-evolution equation for the magnetization [2]
@@
the transition rate when $K=0$ in the presence of a magnetic field $h$. The steady-state global order for $K=0$, denoted as $m^{\mu}_{0,\rm s}$, reads [5]
@@
where it follows from a Taylor expansion of Eq. (S3) that
@@
S6.4 Bound on perturbations of global order
@@
This is our first main result. We are left with determining an upper bound for $\lim_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, for which we use the following theorem shown in [6] (Theorem 2.1):
@@
Since it is known that the finite volume Glauber dynamics on $\mathbb{Z}^{d}$ is exponentially weakly ergodic (see Theorem 3.3 on page 117 in [7] together with proposition 3.9 on page 124 or simply use Eq. (3.15) in [7]), we can directly use Theorem 1. To translate the results from Theorem 1 into a bound for $|\delta m^{\mu}_{\rm s}|$, we first want to bound the subordinate norm for the perturbed infinitesimal generator. To do so, note that the perturbation to the transition rates obeys the following bound,
@@
Note that our proof does not state anything about the regime with coherent steady-state oscillations on the square lattice. Similar to the results in [1], we also observe that spiral defects can destroy coherent steady-state oscillations on the square lattice. This, however, does not have any implication for the regime with static order, which is our main focus in this Section.
S6.6 What happens in the thermodynamic limit?
What happens to the static global order when we take $N\rightarrow\infty$? It is known that the infinite volume Glauber dynamics is not ergodic as the Gibbs measure is not unique [7]. Therefore, one cannot use Theorem 1. However, in the presence of a magnetic field with $h>0$ there exists a region $J\geq J_{0}$ (this includes the infinite $J\rightarrow\infty$ limit) such that the infinite volume Glauber dynamics has exponential convergence, i.e., is weakly ergodic, as shown in Theorem 5.1b in [8] (page 479). Hence, in this regime we can use Theorem 1 to obtain a bound for $|K|\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, which would also be given by Eq. (S108) (albeit that $b$ and $c$ might change due to the presence of $h>0$). Looking at Eq. (S110) we can then take a scaling limit $N\rightarrow\infty$ and $|K|N={\rm constant}$, such that the perturbed steady-state long-range order can get arbitrarily close to the unperturbed value by simply changing the proportionality constant. Hence, we conclude that there exists a regime for $|K|N={\rm constant}$ where the static order is preserved when
@@
Appendix S7 Kinetic Monte-Carlo simulations
For the results shown in Fig. 1c,d (black dashed lines) and Fig. 3 in the Letter we performed kinetic MC simulations on three different types of lattices: (i) the fully-connected mean field (MF) lattice, (ii) the Bethe lattice, and (iii) the square lattice with periodic boundary conditions. Simulations on the Bethe lattice were performed using the random graph algorithm [9, 10], which works as follows: Let us consider a Bethe lattice with coordination number $z$. First, we create a Cayley tree of $i=\{1,...,N\}$ spins with coordination number $z$. The spins on the outer layer are only connected to one spin on the inner layer. To create the remaining $z-1$ connections, we randomly pair spins on the outer layer to other spins on the outer layer. The final result is a Cayley tree with random connections on the outer layer. Note that for both lattices $a$ and $b$ we create new random connections. For large $N$ it has been shown that the Ising model on an ensemble of such random graphs is equivalent to the Ising model on a Bethe lattice [9]. Indeed, for large $N$ we find perfect agreement between the simulations and our theory, as shown in Fig. 1c,d in the main Letter. For the MF lattice, we connect all spins with each other, resulting in a fully connected graph.
Figure S2: Oscillations in the global order $m^{\mu}(t)$ on the mean-field lattice (a), Bethe lattice (b), and square lattice (c), obtained by Monte-Carlo simulations. In each panel, we plot 10 independent trajectories for the magnetization on lattice $a$ (green) and $b$ (blue). Fluctuations between the trajectories lead to small differences over time, which are decreasing with increasing system size, a signature of self-averaging. The values for $J$ and $K$ for each of the lattices are reported in table S1.
S7.1 Simulation setup
In Table S1 we summarize the system size, number of trajectories, and parameter settings, which were used to obtain the spectral density shown in Fig. 3 in the Letter. As initial conditions, we selected a randomly mixed configuration of up and down spins for fixed magnetization. A subset of trajectories for the global order are displayed in Fig. S2, and in each of the lattices we observe coherent oscillations.
Table S1: Simulation parameters for results shown in Fig. 3 in the main Letter.
lattice size ($N$) MC steps # trajectories $J$ $[k_{\rm B}T]$ $K$ $[k_{\rm B}T]$
mean field $1500$ $1500\times 10^{3}$ $500$ $1.5/N$ $0.3$
Bethe $1457$ $1457\times 10^{3}$ $500$ $0.5$ $0.3$
square $40\times 40$ $1600\times 10^{3}$ $500$ $0.6$ $0.3$
To obtain the results shown in Fig. 1c,d we used a Bethe lattice with system size $N=118097$. Such a large system size was not feasible for the setup of Fig. 3 since the spectral density $\langle|\hat{\Psi}^{\mu}_{k}(\omega)|\rangle$ must be averaged over many independent trajectories, resulting in memory issues for too large $N$.
S7.2 Videos
To visualize the dynamics of the nonreciprocal Ising model on the mean-field lattice, Bethe lattice, and square lattice, we have provided videos as Supplementary Material. In each of the videos, we show three independent simulations on the aforementioned lattices, with the coupling strengths $(J,K)$ reported in Table S1. Further information about the videos is given below:
  • In the video “MF_lattice.gif” simulations on the mean-field lattice are shown for $N=2000$ spins, where each spin is connected to every other spin. For illustrative purposes, the edges between spins are not shown.
  • In the video “Bethe_lattice.gif” simulations on the Bethe lattice are shown for $N=13121$ spins, corresponding to a Bethe lattice with $8$ layers and a coordination number of $z=4$. For illustrative purposes, only the first $6$ layers of the Bethe lattice are shown.
  • In the video “Square_lattice.gif” simulations on the square lattice are shown for $N=122\times 122$ spins.
S7.3 Eigenvalues of Laplacian matrix
Figure S3: Eigenvalues $l_{k}$ of the Laplacian matrix $\mathbf{L}$ for the mean-field lattice (a), Bethe lattice (b), and square lattice (c). For the mean-field lattice, all eigenvalues except for the first are degenerate with value $l_{2}=...=l_{N}=N$, where $N$ is the system size given in table S1. The Bethe lattice has a spectral gap between the lowest eigenvalue $l_{1}$ and $l_{2}$.
In Fig. 3 in the main Letter we plot the spectral density $\langle|\hat{\Psi}^{\mu}_{k}(\omega)|\rangle$ as a function of the eigenvalues $l_{k}$ of the Laplacian matrix $\mathbf{L}$. To obtain the eigenvalues, we numerically diagonalized the Laplacian $\mathbf{L}$ in Python, and the resulting eigenvalues are shown in Fig. S3. Note that for the mean-field lattices all eigenvalues except for the first are degenerate and equal to $l_{2}=..=l_{N}=N$.

### v3

@@ Appendix G Kinetic Monte-Carlo simulations
Here, we investigate the stability of steady-states of Eqs. (8)-(10) in the main manuscript, in combination with monomer and Bethe-Guggenheim (BG) approximation, using linear stability analysis. This allows us to identify the region in parameter space where we have a so-called Hopf bifurcation, which marks the transition from a (non-oscillatory) steady-state to coherent oscillations.
S2.1 The “monomer” approximation
Since in the monomer approximation the local order is slaved by the global order, it suffices to consider the linear stability analysis for the global order. Let us consider a small perturbation of the global order around the steady-state value $\textbf{m}_{\rm s}$, i.e., $\mathbf{m}(t)=\mathbf{m}_{\rm s}+\mathbf{\delta m}(t)$. The most interesting steady-state value to consider is given by the disordered state $\mathbf{m}_{\rm s}=0$, in which case the probability (S6)-(S7) can be expanded as
@@
Inserting this linearized expression back into Eq. (8) of the main manuscript, we eventually obtain
@@
Note that $\hat{M}_{\mu\mu}$ can be written in terms of the steady-state solution for the local order, i.e., $\hat{M}_{\mu\mu}=zq^{\mu\mu}_{s}-1$, which we will make use of in the remaining calculation. Furthermore, the off-diagonals are related to the steady-state of the local order between the lattices, i.e., $\hat{M}_{ab}+\hat{M}_{ba}=2q^{ab}_{s}$. Hence, a perturbation of the global order couples to the steady-state value of the local order. The solution of Eq. (S12) can be written in terms of an eigenmode expansion
@@
Since all matrix entries in (S12) are real, the characteristic polynomial also has real coefficients. Therefore, if the eigenvalues are complex, they come in complex-conjugate pairs. The perturbation $\mathbf{\delta m}(t)$ grows in time when ${\rm Re}(\hat{\lambda}_{\pm})>0$, and shrinks in time when ${\rm Re}(\hat{\lambda}_{\pm})<0$. The imaginary part of the eigenvalues tells us whether the perturbation develops oscillations in time ${\rm Im}(\hat{\lambda}_{\pm})\neq 0$, or is monotonic in time ${\rm Im}(\hat{\lambda}_{\pm})=0$. The Hopf bifurcation, also known as an oscillatory instability or type-${\rm II}_{\rm o}$ instability in the Cross-Hohenberg classification Cross and Hohenberg (1993), occurs when the complex conjugate eigenvalues cross the imaginary axis in the complex plane. Based on Eq. (S17), this occurs when
@@
S2.1.1 Perfectly nonreciprocal setting
@@
where the latter equation is directly satisfied for $K\neq 0$. Hence, also in the “monomer” expression we find a critical value for the local order, which is given by $q_{\rm crit}\equiv 1/z$.
S2.2 The Bethe-Guggenheim “pair” approximation
@@
For the sake of simplicity, we directly focus on the perfectly nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$. Upon inserting the steady-state values, the probability can be expanded up to first order in $\delta\mathbf{m}(t)$ and $\delta\mathbf{q}(t)$, which reduces Eqs. (8)-(10) in the main paper to the following linear set of equations
@@
where $\mathbf{M}(q_{s};J,K)$ is a $2\times 2$ matrix and $\mathbf{Q}(q_{s};J,K)$ a $3\times 3$ matrix. Due to the diagonal block structure of the linearized equations, we can handle the perturbations for $\delta\mathbf{m}(t)$ and $\delta\mathbf{q}(t)$ separately. The linear stability of $\delta\mathbf{m}(t)$ is already discussed in the main paper, and here we proceed with $\delta\mathbf{q}(t)$. The elements of the $3\times 3$ matrix $\mathbf{Q}$ are given by
@@
where for the last inequality we again used the triangle inequality. This establishes that $Q_{11}\leq 0$ and $Q_{33}\leq 0$, and therefore when $\tilde{\lambda}_{2,3}$ become complex, their real part obeys the bound ${\rm Re}(\tilde{\lambda}_{2,3})=(Q_{11}+Q_{33})/2\leq 0$. Hence, up to first order, any perturbation $\delta\mathbf{q}(t)$ decays over time. The coherent oscillations in $\mathbf{q}(t)$ observed in Fig. 1c,d in the main paper are therefore an inherently nonlinear effect related to the coupling between $\mathbf{m}(t)$ and $\mathbf{q}(t)$.
@@
S3.1 Recap of the global order
@@
where $\{\boldsymbol{\sigma}\}$ denotes the set of all possible spin configurations $\boldsymbol{\sigma}$. From the master equation (see Eq. (2) in the main paper) we can obtain the time-evolution equation for the magnetization Glauber (1963)
@@
the transition rate when $K=0$ in the presence of a magnetic field $h$. The steady-state global order for $K=0$, denoted as $m^{\mu}_{0,\rm s}$, reads Yang (1952)
@@
where it follows from a Taylor expansion of Eq. (3) in the main paper that
@@
S3.4 Bound on perturbations of global order
@@
This is our first main result. We are left with determining an upper bound for $\lim_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, for which we use the following theorem shown in Mitrophanov (2003) (Theorem 2.1):
@@
Since it is known that the finite volume Glauber dynamics on $\mathbb{Z}^{d}$ is exponentially weakly ergodic (see Theorem 3.3 on page 117 in Bertoin et al. (2004) together with proposition 3.9 on page 124 or simply use Eq. (3.15) in Bertoin et al. (2004)), we can directly use Theorem 1. To translate the results from Theorem 1 into a bound for $|\delta m^{\mu}_{\rm s}|$, we first want to bound the subordinate norm for the perturbed infinitesimal generator. To do so, note that the perturbation to the transition rates obeys the following bound,
@@
Note that our proof does not state anything about the regime with coherent steady-state oscillations on the square lattice. Similar to the results in Avni et al. (2023), we also observe that spiral defects can destroy coherent steady-state oscillations on the square lattice. This, however, does not have any implication for the regime with static order, which is our main focus in this Section.
S3.6 What happens in the thermodynamic limit?
What happens to the static global order when we take $N\rightarrow\infty$? It is known that the infinite volume Glauber dynamics is not ergodic as the Gibbs measure is not unique Bertoin et al. (2004). Therefore, one cannot use Theorem 1. However, in the presence of a magnetic field with $h>0$ there exists a region $J\geq J_{0}$ (this includes the infinite $J\rightarrow\infty$ limit) such that the infinite volume Glauber dynamics has exponential convergence, i.e., is weakly ergodic, as shown in Theorem 5.1b in Martinelli and Olivieri (1994) (page 479). Hence, in this regime we can use Theorem 1 to obtain a bound for $|K|\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, which would also be given by Eq. (S58) (albeit that $b$ and $c$ might change due to the presence of $h>0$). Looking at Eq. (S60) we can then take a scaling limit $N\rightarrow\infty$ and $|K|N={\rm constant}$, such that the perturbed steady-state long-range order can get arbitrarily close to the unperturbed value by simply changing the proportionality constant. Hence, we conclude that there exists a regime for $|K|N={\rm constant}$ where the static order is preserved when